OpenStudy (anonymous):

@Ahaanomegas what is an equation for the line with slope 2/3 and y-intercept 9? a. y=2/3x b. y=9x c. y=2/3x+9 d. y=9x+2/3 what is the equation in slope intercept form for the line that passes through the points (1,-3) and (3,1)? a. y=3x+1 b. y=x-3 c. y=2x+5 d. y=2x-5

OpenStudy (anonymous):

#1: C, by definition of slope-intercept form. #2: The slope is \[ \frac {1 - (-3)}{3 - 1} = \frac {4}{2} = 2 \]and it crosses the point (1, -3), so the y-intercept is b where -3 = 1*2 + b -> b = -5, so y = 2x - 5. (D)